A 21' by 21' square swimming pool is surrounded by a path of uniform width. if the area of the path is 184 sq. ft. find the width of the path.
Answer by wuwei96815(245)

Let x = width of unknown perimeter

Area of path + area of pool = x^2
442 + 184 = x^2
626 = x^2
25 = x

Path = difference between outer perimeter and pool perimeter = 25 - 21 = 4 ft.

Does that look right?